OpenStudy (cutiecomittee123):

tan(25pi/2)= How do i solve for this? The options are: 0 1 -1 undefined

OpenStudy (javalos01):

hello do you need help

OpenStudy (cutiecomittee123):

yes. but i already found out the answer to this question here is another that i need help with if you are willing to help that would be awesome "The statement tantheta=-12/5 and csctheta==13/12 and the terminal point determined by theta is in quadrant 3....... a.cannot be true because 12^2+5^2 does not equal 1" b.cannot be true because if tantheta=-12/5 then csctheat=+-13/5" c.cannot be true because tantheta must be less than one" d.cannot be true because tantheta is greater than zero in quadrant 3" Which one of these fits the best and why?

OpenStudy (cutiecomittee123):

@javalos01

OpenStudy (cutiecomittee123):

I know that if it is in quadrant 3 then sintheta is negative and costheta is positive.

OpenStudy (javalos01):

ok what do u need

OpenStudy (javalos01):

correct go on

OpenStudy (javalos01):

if all of them are wrong there has to be a correct one

OpenStudy (javalos01):

which one do you for sure think it is

OpenStudy (cutiecomittee123):

i think that a is the correct one because when you are trying to solve for a trigonometric eqtuation and you have two sides you always set it equal to one. and if 12^2+5^2 arent equal to 1 it cannot be true

OpenStudy (javalos01):

ok i gtg and i think ur correct

OpenStudy (cutiecomittee123):

well thanks

OpenStudy (anonymous):

\[\tan \frac{ 25 \pi }{ 2 }=\tan \left( 6\times 2 \pi+\frac{ \pi }{ 2 } \right)=\tan \frac{ \pi }{ 2 }=in finity\]

OpenStudy (anonymous):

infinity means undefined.

OpenStudy (cutiecomittee123):

oh i see thanks for that!

OpenStudy (cutiecomittee123):

can you help with my other question. did i chose the right answer?

OpenStudy (anonymous):

sure ,if possible.

OpenStudy (cutiecomittee123):

alright its on here already and that guy tried to help me but iw as wondering if you can double check it for me its the question above ^^^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your statement is wrong tan theta is positive in quadrant 3

OpenStudy (anonymous):

here csc theta is positive, tan theta is negative so theta lies in 2nd quadrant.

OpenStudy (anonymous):

d is correct.

OpenStudy (anonymous):

d says statement is wrong because tan theta >0 in quadrant 3.

OpenStudy (cutiecomittee123):

but how would tantheta be postitive when its y/x and y is negative -y/x= a negative outcome

OpenStudy (cutiecomittee123):

@surjithayer

OpenStudy (anonymous):



OpenStudy (cutiecomittee123):

oh yeah. okay gotcha! and that cannot work so it must be in the second quadrant because tantheta is greater than 0 in quadrant 3 and in the statment it says that tan is negative

OpenStudy (cutiecomittee123):

awesome thanks

OpenStudy (anonymous):

yw